OpenStudy (sheriph05):

Guys, please help with this question. Given that S = {S_n;S_n = 1+\sum_{i=1}^{n^2} (\frac{1}{i!}), n = 1,2... } Show that 3\in U_s

OpenStudy (mathmale):

I see you've tried to present this problem properly using the Equation Editor, but for some reason did not succeed. Could you possibly draw this equation in the Draw utility, below?

satellite73 (satellite73):

\[ S = {S_n;S_n = 1+\sum_{i=1}^{n^2} (\frac{1}{i!}), n = 1,2... } \]

satellite73 (satellite73):

Show that \(3\in U_s\)

satellite73 (satellite73):

some confusion here clearly there is no \(U_S\) in the question

OpenStudy (sheriph05):

I am confused on this as well as i cannot identify what U_s stands for

ganeshie8 (ganeshie8):

I think they want you show \(3 \not\in S\)

OpenStudy (zzr0ck3r):

Some form of union?

OpenStudy (holsteremission):

Looks like \(S_n=\exp_{n^2}(1)=\displaystyle\frac{e\Gamma(n^2+1,1)}{\Gamma(n^2+1)}\) where \(\exp_n(x)=\displaystyle\sum_{i=0}^n\frac{x^i}{i!}\) denotes the exponential sum function http://mathworld.wolfram.com/ExponentialSumFunction.html and \(\Gamma(n,x)\) and \(\Gamma(n)\) the (incomplete) gamma functions http://mathworld.wolfram.com/IncompleteGammaFunction.html http://mathworld.wolfram.com/GammaFunction.html Not that that gives any more information... Are you sure you've never seen this notation defined or used anywhere else?

OpenStudy (sheriph05):

It is about showing that 3 is an element of upper bound of the set S

ganeshie8 (ganeshie8):

\[\begin{align}1+\color{blue}{\sum_{i=1}^{\infty} (\frac{1}{i!})} &= 1+\color{blue}{\dfrac{1}{1} + \dfrac{1}{1\cdot 2} + \dfrac{1}{1\cdot 2\cdot 3} + \cdots}\\~\\ &\lt 1 + \color{blue}{\dfrac{1}{1} + \dfrac{1}{2} + \dfrac{1}{2\cdot 2} +\dfrac{1}{2\cdot 2\cdot 2}+ \cdots}\\~\\ &= 1+ \color{blue}{2} \end{align}\] Since the sequence \(S_n\) is increasing, all the terms in the sequence are less than 3.

OpenStudy (sheriph05):

@ganeshie8 Okay..Thanks for the insight....I get the 1st Line as this appears to be order of factorials...However i am still in dark as regards the second line.... can u please explain the second line?

ganeshie8 (ganeshie8):

Notice that \(1\cdot 2 \cdot 3 \gt 2\cdot 2\)

ganeshie8 (ganeshie8):

That means \(\dfrac{1}{1\cdot 2 \cdot 3} \lt \dfrac{1}{2\cdot 2}\)

ganeshie8 (ganeshie8):

\[\begin{align}1+\color{blue}{\sum_{i=1}^{\infty} (\frac{1}{i!})} &= 1+\color{blue}{\dfrac{1}{1} + \dfrac{1}{1\cdot 2} + \dfrac{1}{1\cdot 2\cdot 3} + \dfrac{1}{1\cdot 2\cdot 3\cdot 4} + \cdots}\\~\\ &\lt 1 + \color{blue}{\dfrac{1}{1} + \dfrac{1}{2} + \dfrac{1}{2\cdot 2} +\dfrac{1}{2\cdot 2\cdot 2}+ \cdots}\\~\\ &= 1+ \color{blue}{2} \end{align}\] Since the sequence \(S_n\) is increasing, all the terms in the sequence are less than 3.

OpenStudy (sheriph05):

thanks